QUESTION 1

A firm uses a single input labour to produce output q according to the production

𝑞 = 8√𝐿

The commodity sells for R150 per unit and the wage rtae is R75 per hour.

a) Find the profit maximizing quantity of L (3)

b) Find the profit maximizing quantity of q and the level of maximum profit. (5)

c) Suppose now the firm is taxed at R30 per unit of output and that the wage rate is subsidized

at a rate of R15 per hour. Assume that the firm is a price taker, so the price of the producer

remains at R150. Find the new profit maximizing levels of L, q and profit. (6)

d) Nowsupposethefirmisrequiredtopaya20percenttaxonitsprofits,Findthenewprofit maximizing levels of L, q and profit.


1
Expert's answer
2021-05-24T13:33:20-0400

Solution:

a.). The profit maximizing quantity of L: This is where the firm’s marginal revenue product of labor (MRP) equals the wage (W).

"MRP_{L} = W"

"MRP_{L} = MPL\times P"

"P = Price \;of\;the\;commodity = 150"

"W = 75"

"MPL = \frac{\partial Q} {\partial L} = \frac{4}{L^{1/2} }"


"Therefore:"

"MRP_{L} = W"

"(\frac{4}{L^{1/2} })\times 150 = 75"


"\frac{600}{L^{1/2} } = 75"


"75L^{1\/2} = 600"


"L^{1/2} = \frac{600}{75} =8"

"Square\; both \;sides:"

"L = 64"


"The \;quantity\; of \;labor \;that \;will \;maximize\; profit = 64"


b.). Profit maximizing quantity of q and the level of maximum profit:

"Q = 8\sqrt{L}"


"Substitute\; L \;in\; the \;equation:"

"Q = 8\sqrt{64}"


"Q = 8\times8 = 64"


"Profit \;maximizing\; quantity\; of\; q = 64"


"The \;level\; of\; maximum\; profit = Price\times Quantity"


"Profit = 150\times64 = 9,600"


"The\; level\; of\; maximum\; profit = R9,600"


c.). "MRP_{L} = W"

"MRP_{L} = MPL\times P"

"P = Price \;of\;the\;commodity = 150"

"W = 75 - 15 = 60"

"MPL = \frac{\partial Q} {\partial L} = \frac{4}{L^{1/2} }"


"Therefore:"

"MRP_{L} = W"

"(\frac{4}{L^{1/2} })\times 150 = 60"


"\frac{600}{L^{1/2} } = 60"


"60L^{1\/2} = 600"


"L^{1/2} = \frac{600}{60} =10"

"Square\; both \;sides:"

"L = 100"


"The \; new\;quantity\; of \;labor \;that \;will \;maximize\; profit = 100"

To get the new profit maximizing quantity of q and the level of maximum profit:

"Q = 8\sqrt{L}"


"Substitute\; L \;in\; the \;equation:"

"Q = 8\sqrt{100}"


"Q = 8\times10 = 80"


"Profit \;maximizing\; quantity\; of\; q = 80"


"The \;level\; of\; maximum\; profit = Price\times Quantity"


"Profit = 150\times80 = 12,000"


"The\; level\; of\; maximum\; profit = R12,000"


d.). "New\;price = 150 + (20 \% \;tax\;increase) = 180"

"MRP_{L} = W"

"MRP_{L} = MPL\times P"

"P = Price \;of\;the\;commodity = 180"

"W = 75 - 15 = 60"

"MPL = \frac{\partial Q} {\partial L} = \frac{4}{L^{1/2} }"


"Therefore:"

"MRP_{L} = W"

"(\frac{4}{L^{1/2} })\times 180 = 60"


"\frac{720}{L^{1/2} } = 60"


"60L^{1\/2} = 720"


"L^{1/2} = \frac{720}{60} =12"

"Square\; both \;sides:"

"L = 144"


"The \; new\;quantity\; of \;labor \;that \;will \;maximize\; profit = 144"

To get the new profit maximizing quantity of q and the level of maximum profit:

"Q = 8\sqrt{L}"


"Substitute\; L \;in\; the \;equation:"

"Q = 8\sqrt{144}"


"Q = 8\times12 = 96"


"Profit \;maximizing\; quantity\; of\; q = 96"


"The \;level\; of\; maximum\; profit = Price\times Quantity"


"Profit = 180\times96 = 17,280"


"The\; level\; of\; maximum\; profit = R17,280"